C=100 + .8(Q - T)

I = 200

G=400

T=100

EX = 300

IM = 400

1.) Please find Household Consumption and Saving

2.) The spending multiplier for this economy is?

3.) To move this economy to Q = 2900 the Government would increase G by:
Answer #1

Answer

The equilibrium is at (Q=GDP=Y)

Q=C+I+G+NX

NX=EX-IM=300-400=-100

Q=100+0.8(Q-100)+200+400-100

Q=600+0.8Q-80

0.2Q=520

Q=2600

1)

consumption =100+0.8*(Q-T)=100+0.8*(2600-100)

=2100

Saving =Q-C-G=2600-2100-400=100

======

2)

Spending multiplier =1/(1-MPC)

MPC=coefficient of (Q-T) in the consumption function=0.8

Spending multiplier =1/(1-0.8)=5

3)

change in GDP=2900-2600=300

required change in G=change in GDP/spending multiplier

=300/5

=60

to move the economy to 2900 the govenment should increase G by 60 units